# A unified framework for equation discovery and dynamic prediction of hysteretic systems

### v1

@@ 5 Conclusion
This research presented a unified framework for equation discovery and dynamic prediction of hysteretic systems, with a proposed solver-based internal variable learning scheme that addresses one of the key challenges in hysteresis modeling – the difficulty of inferring internal variables. We summarize existing work on equation discovery for hysteretic systems and categorize them according to the level of assumed knowledge available. The proposed unified framework in our work is capable of addressing two major challenges of hysteretic systems across various scenarios: the learning of unobservable internal variables and the discovery of governing equations with complex nonlinear expressions. The effectiveness of the framework is validated through three representative case studies. On the Bouc-Wen benchmark, it accurately recovers the governing equations even in the presence of noise, confirming its robustness and reliability. For a more complex nonlinear structure with cubic stiffness and fractional hysteretic exponents, the framework successfully achieves superior accuracy and interpretability compared with library-based methods such as SINDy. In the real shake-table experiment of a yielding structure, the framework reveales concise rate-dominated dynamics without NSD and more complex nonlinear hysteresis behavior with NSD, achieving stable predictions.
This research not only provides insights into equation discovery and dynamic prediction of hysteretic systems but also demonstrates the potential of the proposed framework for broader applications in nonlinear dynamical systems, including those with or without internal variables. Its flexibility and interpretability make it applicable to various domains such as structural dynamics, material modeling, and control system identification, where nonlinear and memory-dependent behaviors are common. Future work will focus on enhancing the efficiency and scalability of the framework, improving its ability to handle higher-dimensional systems, and exploring probabilistic or hybrid modeling strategies to further extend its applicability to real-world engineering and scientific problems.
@@ Acknowledgments
The authors would like to express their sincere gratitude to Prof. Satish Nagarajaiah at Rice University for generously sharing the experimental data of SDOF yielding structure that made this research possible. The authors wish to express their gratitude for the financial support received from Guangdong Provincial Fund - Special Innovation Project (2024KTSCX038); Research Grants Council of Hong Kong through the Research Impact Fund (R5006-23); the Guangdong Provincial Key Lab of Integrated Communication, Sensing and Computation for Ubiquitous Internet of Things (No. 2023B1212010007).

### v2

@@ 5 Conclusion
This research presented a unified framework for equation discovery and dynamic prediction of hysteretic systems, with a proposed solver-based hysteretic variable learning scheme that addresses one of the key challenges in hysteresis modeling – the difficulty of inferring hysteretic variables. We summarize existing work on equation discovery for hysteretic systems and categorize them according to the level of assumed knowledge available. The proposed unified framework in our work is capable of addressing two major challenges of hysteretic systems across various scenarios: the learning of unobservable hysteretic variables and the discovery of governing equations with complex nonlinear expressions. The effectiveness of the framework is validated through representative case studies. On the Bouc-Wen benchmark, it accurately recovers the governing equations even in the presence of noise, confirming its robustness and reliability. For a more complex nonlinear structure with cubic stiffness and fractional hysteretic exponents, the framework successfully achieves superior accuracy and interpretability compared with library-based methods such as SINDy . In the real shake-table experiment of a yielding structure, the framework reveals concise rate-dominated dynamics without NSD and more complex nonlinear hysteresis behavior with NSD, achieving stable predictions.
This research not only provides insights into equation discovery and dynamic prediction of hysteretic systems but also demonstrates the potential of the proposed framework for broader applications in nonlinear dynamical systems, including those with or without hysteretic variables. Its flexibility and interpretability make it applicable to various domains such as structural dynamics, material modeling, and control system identification, where nonlinear and memory-dependent behaviors are common. Future work will focus on enhancing the efficiency and scalability of the framework, improving its ability to handle more complex systems, and exploring probabilistic or hybrid modeling strategies to further extend its applicability to real-world engineering and scientific problems.
@@ Acknowledgments
The authors would like to express their sincere gratitude to Prof. Satish Nagarajaiah at Rice University for generously sharing the experimental SDOF yielding structure data that made this research possible. The authors wish to express their gratitude for the financial support received from Guangdong Provincial Fund - Special Innovation Project (2024KTSCX038); Research Grants Council of Hong Kong through the Research Impact Fund (R5006-23); Guangdong Provincial Key Lab of Integrated Communication, Sensing and Computation for Ubiquitous Internet of Things (No. 2023B1212010007); Natural Science Foundation of Guangdong Province, China (General Fund; Grant No. 2026A1515011629).
Appendix A Network architecture and training details
In this work implementation, $\mathcal{N}_{\theta}(\cdot)$ in Full Equation Discovery is parameterized as a multilayer perceptron with two hidden layers, each of width 128, and hyperbolic tangent activation functions. The network parameters are optimized jointly within the solver-based training procedure using the Adam algorithm, with a learning rate of $3\times 10^{-4}$, first- and second-moment decay rates of 0.9 and 0.98, respectively, and an $\ell_{2}$ regularization coefficient of $10^{-4}$. To improve optimization stability, gradient norm clipping with a threshold of 1.0 is applied. Training is terminated either upon convergence of the loss or when a prescribed maximum number of iterations (e.g., 10,000) is reached.
Appendix B Computational complexity and scalability
We briefly discuss computational complexity and scalability as the system dimension increases. Let $r$ denote the number of degrees of freedom (DOFs), $q$ the number of hysteretic variables, and $N$ the total number of sampled time points.
B.1 Step 1: Solver-based hysteretic variable learning.
At this step, the trainable hysteretic variable is learned by embedding the governing dynamics into a differentiable ODE solver and jointly optimizing the model parameters and the trainable hysteretic variables through trajectory matching. Since the trainable hysteretic variable is parameterized outside the solver (shown in Figure 2), the embedded solver advances only the displacement and velocity states, so the solver state has dimension $2r$. For each optimization iteration, the solver is unrolled over $N$ time samples, and the trainable hysteretic variable is injected into the motion equation at the corresponding sampled times. If the solver uses $n_{\mathrm{s}}$ stage evaluations per time step, with $n_{\mathrm{s}}=4$ for RK4, the per-iteration solver cost is $\mathcal{O}(Nn_{\mathrm{s}}\,2r)$, while updating the trainable hysteretic variables over all time samples introduces an additional cost of $\mathcal{O}(Nq)$ per iteration. Hence, the per-iteration computational cost is $\mathcal{O}\!\left(Nn_{\mathrm{s}}\,2r+Nq\right)$. Since $n_{\mathrm{s}}$ is typically small and fixed in practice, this can be simplified to:
$\mathcal{O}\!\left(N(2r+q)\right).$ (25)
B.2 Step 2: Equation discovery.
At this step, SR is applied to the learned trajectories to recover explicit governing equations. The computational cost is dominated by the symbolic search process. Let $P$ denote the population size and $G$ the number of generations. For each candidate symbolic expression, the main cost comes from evaluating the expression over the available data samples. If $C$ denotes the per-sample cost of evaluating one candidate symbolic expression, then the computational cost of Step 2 is:
$\mathcal{O}\!\left(N\,P\,G\,C\right).$ (26)
Relative to conventional gradient-based parameter estimation for a prescribed hysteretic model, the proposed Step 1 shares a similar computational backbone, as both require repeated forward ODE integration and gradient-based optimization over trajectory data. However, the present framework is generally more computationally demanding in practice because the hysteretic relation is not fixed a priori and must instead be learned in a more flexible function space, rather than calibrated through a small set of prescribed physical parameters. In addition, the proposed framework introduces an offline Step 2 after training. These additional computational costs represent a trade-off for reduced structural assumptions and the ability to recover an interpretable symbolic governing equation.
B.3 Impact of multiple hysteretic variables.
When hysteresis is present at multiple or adjacent DOFs, different hysteretic variables may contribute to overlapping restoring-force components because of structural coupling. As a result, similar observable responses can be explained by different combinations of hysteretic variables, which makes the different hysteretic contributions harder to identify and renders the joint optimization more ill-conditioned. From both computational and identifiability perspectives, scalability therefore relies not only on computational resources, but also on sufficiently informative excitations, multiple training trajectories, adequate sensing, and physically motivated regularization (e.g., sparsity or locality in cross-DOF coupling) to stabilize learning and improve the separability of hysteretic contributions.
Appendix C Additional experiments on the Bouc-Wen hysteretic system benchmark
To further validate the effectiveness and practical applicability of the proposed framework, we conduct some additional benchmark experiments beyond the baseline setting.
C.1 Cross-excitation validation
To assess whether the discovered equations are biased to a specific excitation history, we swapped the training and testing roles of the two excitation signals used in the baseline in the medium-noise setting (SNR = 20 dB). Specifically, the multisine signal is used for training and the sinesweep signal for testing, i.e., the reverse of the setting shown in Figure 3. For a fair comparison, all remaining training settings (data length, sampling rate, optimization hyperparameters, etc.) are kept identical to the baseline.
For quantitative evaluation, Table LABEL:tab:cross&extra_exp1 reports the NRMSE results, together with discovered equations $1.7803\ddot{x}+8.0472\dot{x}+47360.7258x+0.8736z=u$ and $\dot{z}=57072.5992\dot{x}-748.504|\dot{x}|z+748.504\dot{x}|z|$. Figures 19 and 20 visualize the predicted responses against the ground truth in the swapped excitation setting. In this cross-excitation validation example, the proposed framework maintains competitive accuracy for both seen and unseen excitation types, suggesting that the discovered equations are not related to a specific input history and retain predictive capability across the two benchmark excitation types.
Figure 19: Displacement results on the benchmark system - cross-excitation validation. In contrast to Figure 5, the multisine signal is used as the training input, while the sinesweep signal is used as the testing input. (a) Training results. (b) Testing 1 (same excitation) results. (c) Testing 2 (unseen excitation) results.
Figure 20: Velocity results on the benchmark system - cross-excitation validation. In contrast to Figure 6, the multisine signal is used as the training input, while the sinesweep signal is used as the testing input. (a) Training results. (b) Testing 1 (same excitation) results. (c) Testing 2 (unseen excitation) results.
C.2 Extrapolation test on unseen amplified excitation
To evaluate performance beyond the training operating range, we keep the same medium-noise baseline setting and test the discovered equations under an amplitude-scaled excitation without retraining. In contrast to the nominal testing condition, we construct an amplitude-extrapolated test input by scaling the multisine excitation by a factor of 1.5, while keeping all other settings unchanged. Without retraining or re-identification, we evaluate the governing equations obtained under the baseline setting.
The quantitative NRMSE results are summarized in Table LABEL:tab:cross&extra_exp1, and the corresponding displacement and velocity comparisons are visualized in Figures. 21 and 22. From this example, the proposed framework remains stable under the increased excitation amplitude and achieves accuracy comparable to that under the nominal testing condition, indicating reasonable extrapolation capability within the considered amplitude range.
Figure 21: Displacement results on the benchmark system - extrapolation test. In contrast to Figure 5, the multisine excitation used for testing is amplitude-scaled by a factor of $1.5$. (a) Training results. (b) Testing 1 (same excitation) results. (c) Testing 2 (unseen excitation) results.
Figure 22: Velocity results on the benchmark system - extrapolation test. In contrast to Figure 6, the multisine excitation used for testing is amplitude-scaled by a factor of $1.5$. (a) Training results. (b) Testing 1 (same excitation) results. (c) Testing 2 (unseen excitation) results.
C.3 Sensitivity analyses
To evaluate robustness to modest parameter mismatch, we keep the training procedure unchanged and perturb the true system parameters only at test time by applying a $\pm 5\%$ drift to $c$ or $k$ while holding the other parameter at its nominal value.
The quantitative results are summarized in Table LABEL:tab:sensitivity_exp1 in terms of NRMSE for both displacement and velocity responses. The performance degradation caused by perturbing $k$ is noticeably larger than that induced by perturbing $c$, indicating that the system responses are more sensitive to stiffness variations in this benchmark. A plausible reason is that stiffness directly determines the restoring-force scale and the oscillation characteristics of the system, whereas damping primarily influences the rate of energy dissipation. Even under these small parameter drifts, the framework maintains reasonably low NRMSE values across both testing cases, indicating robust predictive performance in the presence of mild parameter uncertainty.
C.4 Noisy wideband random excitation training
To assess whether the proposed framework is suitable for noisy and new excitation, a band-limited wideband random excitation is synthesized in the frequency domain and transformed back to the time domain, after which measurement noise is added to emulate a noisy input signal. The same medium-noise setting is adopted for both the system responses and the excitation measurement.
Table A unified framework for equation discovery and dynamic prediction of hysteretic systems reports the NRMSE results. In this experiment, the model is trained with an excitation signal different from that in the baseline, while the testing excitation is kept unchanged for a fair comparison. Figures 23 and 24 visualize the predicted displacement and velocity responses. The results show that the proposed framework maintains good predictive accuracy under the new and noisy excitation condition.
Figure 23: Displacement results on the benchmark system - noisy wideband random excitation training. (a) Training results. (b) Testing 1 (same excitation) results. (c) Testing 2 (unseen excitation) results.
Figure 24: Velocity results on the benchmark system - noisy wideband random excitation training. (a) Training results. (b) Testing 1 (same excitation) results. (c) Testing 2 (unseen excitation) results.
C.5 Without hysteresis
To further assess practical applicability, we test whether the framework reduces to an appropriate non-hysteretic model when hysteresis is absent. We construct a non-hysteretic benchmark by setting $\beta=0$ and $\gamma=0$ in Eq.(15), so that the hysteretic link equation reduces to $\dot{z}(t)=A\dot{x}(t)$. Neglecting the integration constant, we obtain $z(t)=Ax(t)$. Therefore, the true Eq.(15) is $2\ddot{x}(t)+10\dot{x}(t)+100000x(t)=u(t)$, degenerate to a non-hysteretic case. All other training and evaluation settings are kept identical to baseline.
In this without hysteresis experiment, the discovered dynamic motion equation is $1.9998\ddot{x}(t)+10.0049\dot{x}(t)+49999.4985x(t)+1.0089z(t)=u(t)$, and the discovered hysteretic link equation is $\dot{z}(t)=49999.5616\dot{x}(t)$. Substituting the hysteretic link relation into the motion equation yields the degenerate non-hysteretic form $1.9998\ddot{x}(t)+10.0049\dot{x}(t)+100444.0562x(t)=u(t)$, which is consistent with the expected non-hysteretic reduction. Figure 25 shows the hysteresis loops. In panel (c), $F=kx+\alpha z=(k+A)x$, and panels (b) and (c) clearly exhibit an approximately linear relationship in the degenerate regime, consistent with the above analysis. The NRMSE results in Table A unified framework for equation discovery and dynamic prediction of hysteretic systems further confirm that the proposed framework remains accurate when the dynamics reduce to the simpler non-hysteretic case. This is also supported by the response comparisons in Figures 26 and 27. These results further support the unified nature of the proposed framework, which can recover both hysteretic and degenerate non-hysteretic dynamics without requiring a prescribed model structure.
Figure 25: Hysteresis loops on the benchmark system - without hysteresis. (a) $x-\dot{x}$ hysteresis loop. (b) xzx-z hysteresis loop. (c) xF(F=(k+A)x)x-F(F=(k+A)x) hysteresis loop.
Figure 26: Displacement results on the benchmark system - without hysteresis. (a) Training results. (b) Testing 1 (same excitation) results. (c) Testing 2 (unseen excitation) results.
Figure 27: Velocity results on the benchmark system - without hysteresis. (a) Training results. (b) Testing 1 (same excitation) results. (c) Testing 2 (unseen excitation) results.